OpenStudy (peterg2001):

help please The ratio of boys to girls is 5 to 4. There are 80 boys. How many girls are there? 4 16 64 100

OpenStudy (lizz123):

ratio of B to G = 5/4 when B = 80 then 5/4 = 80/x cross multiply then 4(80) = 5x 320 = 5x x = 64 answer C